OpenStudy (anonymous):

Please Help!!!!! A farmer has a rectangular garden plot surrounded by 200 ft of fence. Find the length and width of the garden if its area is 2400ft^(2)

OpenStudy (souvik):

perimeter=2(l+w) Area=l*w

OpenStudy (anonymous):

Let the length be l, width be w. From the data, l+b=200 l*b=2400 Solve both, you will get a quadratic eqn. And the answer.

OpenStudy (anonymous):

Sorry it is 2(l+b)

OpenStudy (anonymous):

Wait what?

OpenStudy (anonymous):

The length of fence is the perimeter. 2(l+b) The area of garden is l*b.

OpenStudy (anonymous):

THE garden is surrounded by 200 feet of fence means itz perimeter =200ft and area is 2400 sq. ft (given) p=2(l+b) area=l*b let the length be l and breadth be b Therefore 200=2(l+b) 2400=lb. solve it u l get the answer

OpenStudy (anonymous):

So its two separate equations?

OpenStudy (anonymous):

yep

OpenStudy (souvik):

yes...you have to solve l and b using this two equations..

OpenStudy (anonymous):

I=100 b=100

OpenStudy (anonymous):

no \[2(100+100)\neq200 and 100.100\neq2400\]

OpenStudy (anonymous):

You have to equations:\[\bf P=2l+2w=200 \implies l+w=100\]\[\bf A=lw=2400\]Re-arrange second 'Area' equation and solve for either 'l' or 'w'. Substitute that in to the first equation and solve for the other variable. @angelita321

OpenStudy (anonymous):

L+(2400)=100?

OpenStudy (souvik):

\(l+2400/l=100\)

OpenStudy (souvik):

now cross multiplication..

OpenStudy (anonymous):

is the lenght 60 and width 40

OpenStudy (souvik):

yeah...well done..

OpenStudy (anonymous):

yep itz 60 and 40

OpenStudy (anonymous):

but my dad helped me cause i didn't understand . we just plug in random #'s

OpenStudy (anonymous):

which part did u not understand???? @angelita321

OpenStudy (anonymous):

everything

OpenStudy (anonymous):

sld i explain???

OpenStudy (souvik):

do you know this formulas?

OpenStudy (anonymous):

a=lw and please and thank you

OpenStudy (anonymous):

do u knw the formulas to find the perimeter and area of a rectangle i mean p=2(l+w) and a=l*w @angelita321

OpenStudy (anonymous):

not the perimeter

OpenStudy (anonymous):

ok then let me first explain u the formula for perimeter

OpenStudy (souvik):



OpenStudy (souvik):

the perimeter of rectangle ABCD is AB+BC+CD+DA=l+w+l+w=2(l+w)

OpenStudy (anonymous):

ohhh !!

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

nw ur question says that 'a rectangular garden plot surrounded by 200 ft of fence' in other words we can say that the perimeter of the rectangular garden is 200ft

OpenStudy (anonymous):

aha

OpenStudy (anonymous):

nw perimter=200ft which means 2(l+w)=200=>l+w=100

OpenStudy (anonymous):

i got that part

OpenStudy (anonymous):

ok have u got the area part??

OpenStudy (anonymous):

@angelita321

OpenStudy (anonymous):

i understood the top part

OpenStudy (anonymous):

nw which part have u not understood??

OpenStudy (anonymous):

but what do you mean the area part. since l+w=100

OpenStudy (anonymous):

means????

OpenStudy (anonymous):

@angelita321

OpenStudy (anonymous):

since l+w=100 wouldn't it just be l=11 and w=100??

OpenStudy (anonymous):

l=100*

OpenStudy (anonymous):

no how can it be??? if l =11 w=100 then l+w=111 but it says l+w=100

OpenStudy (anonymous):

okay and what do i do next?

OpenStudy (anonymous):

see area part i think u have undrstd that???

OpenStudy (anonymous):

area = lw

OpenStudy (anonymous):

2400=100l timies 100w

OpenStudy (anonymous):

ok nw we can see that l+w=100 lw=2400

OpenStudy (anonymous):

ahs

OpenStudy (anonymous):

aha*

OpenStudy (anonymous):

nw solve these two equations u l get the value of l and w

OpenStudy (anonymous):

l+w=100 lw=2400 w100=2400 w=24 ???

OpenStudy (anonymous):

no noo thtz nt the wat to slv

OpenStudy (anonymous):

i'm lost. i'll ask tomorrow again. it's late. thanks for helping though